# Radio observations point to a moderately relativistic outflow in the fast X-ray transient EP241021a

### v1

@@ 3.3 X-ray Observations
Observations of EP241021a were conducted with the Follow-up X-ray Telescope (FXT) onboard Einstein Probe, starting $\approx$36.5 hours after the EP-WXT trigger (Wang et al., 2024d). Follow-up observations from the Swift X-ray Telescope (XRT) were reported by Busmann et al. (2025), revealing a long-lasting X-ray counterpart. To determine whether the long-lived nature of the X-ray emission was possibly due to contamination from a nearby source, we analyzed public data from the Chandra X-ray Observatory, which observed EP241021a approximately 14 days after trigger (ObsID: 30598; PI: Jonker) for a total exposure of 9.9 ks. A single X-ray source is detected within the XRT localization at coordinates RA = 01h 55m 23.3984s, Dec = +05 56 17.498′′ (RA = 28.847, Dec = +5.938) with $90\%$ positional uncertainty of $0.52^{\prime\prime}$. Source counts were extracted using the CIAO software package (v4.16) [^2] https://cxc.harvard.edu/ciao/, employing a circular aperture with a radius of 1.1 arcseconds, corresponding to approximately 90% encircled energy of the Chandra point spread function (PSF). Aperture photometry was performed in the 0.5–8.0 keV band. The resulting spectrum was modeled using an absorbed power-law model (tbabs*powerlaw) in XSPEC, with the hydrogen column density ($N_{\mathrm{H}}$) fixed to the Galactic value reported in Hu et al. (2024b). We find a best-fit photon index of $\Gamma=1.3\pm 0.7$, and an unabsorbed model flux of $(6.1^{+1.4}_{-1.2})\times 10^{-14}$ erg cm-2 s-1 in the 0.3–10 keV band. This value is consistent with the nearly simultaneous XRT measurements (Busmann et al., 2025) and confirms that the late-time flattening is an intrinsic property of the source.
@@ 4 Constraints from radio observations
As mentioned in the Section 1, afterglows from compact radio sources are subject to ISS, which can produce significant flux modulations due to distorted wavefronts (Walker, 1998). Specifically, such flux variations can arise from diffractive and refractive scintillation, each sensitive to different angular scales and frequency bandwidths (Narayan, 1992; Goodman, 1997). By identifying the epochs when diffractive scintillation ceases and refractive scintillation becomes suppressed, we can place lower and upper bounds, respectively, on the angular size of the source. These size constraints, when combined with the angular diameter distance to the source, allow us to infer limits on the Lorentz factor of the outflow, as demonstrated in previous studies of GRBs and other relativistic transients (e.g., Frail et al., 2000; Cenko et al., 2008; Zhang et al., 2023). Motivated by this, in Section 4.1, we present evidence for ISS during the radio monitoring of EP241021a. In Section 4.2, we use the ISS behavior to constrain the source’s angular size and Lorentz factor. We calculate the brightness temperature in Section 4.3 using early-time radio data.
@@ 4.1 Evidence for ISS
The strength and timescale of ISS effects depend on both the observing frequency and the angular size of the source (e.g., Liu et al., 2022a). ISS manifests on two distinct timescales: diffractive ISS (DISS), which induces large-amplitude and rapid fluctuations over minutes due to small-scale interference patterns; and refractive ISS (RISS), which causes slower variability over weeks to months as the signal traverses larger-scale density structures in the scattering screen (Walker, 1998; Lewandowski et al., 2011; Granot & van der Horst, 2014; Liu et al., 2022b). These scintillation effects can produce substantial variations in both flux and spectral shape, especially during the early stages of an afterglow. Initially, when the GRB jet is highly-collimated, DISS dominates the variability; as the emitting region expands, DISS becomes suppressed, and RISS emerges as the dominant contributor (Frail et al., 1997). Consequently, different radio facilities, operating at various frequencies and timescales, might detect variability on distinct timescales.
@@
Assuming that scattering occurs in a thin screen along the line of sight, the screen is located at a distance $D_{\text{scr}}\sim(h_{z}/2)\sin^{-1}|b|\approx 0.6$ kpc from the observer (Reynolds, 1989; Galama et al., 2003), where $h_{z}$ is the scale height of the ionised gas layer ($\sim 1$ kpc) and $b=-53.47$ is the Galactic latitude in degrees. We use the NE2001 model of the Galactic distribution of free electrons (Cordes & Lazio, 2002; Walker, 1998) to derive the scattering measure (SM) of $1.3\times 10^{-4}\,\text{kpc}\,\text{m}^{-20/3}$ and a transition frequency of $\nu_{0}\sim 6$ GHz. Below this frequency, density irregularities in the ISM cause strong phase perturbations leading to significant scattering, whereas above this frequency, scattering effects become weaker.
The decorrelation bandwidth ($\Delta\nu_{\rm dc}$) for diffractive scintillation represents the frequency range over which the intensity variations caused by diffraction remain correlated (Goodman, 1997):
@@ 4.2 Angular size measurements and Lorentz factor
where $\theta_{F_{0}}$ denotes the angular size of the first Fresnel zone at the scattering screen, yielding a value of $\theta_{F_{0}}=2.0$ $\mu$as (Granot & van der Horst, 2014) for EP241021a. After the second epoch, we observe no significant short-term variation in the measured flux density, indicating that DISS has ceased by $\approx$18 days post-burst. Thus, the source angular size $\theta_{s}$ must be larger than a critical threshold, $\theta_{d}=\theta_{F_{0}}\times(\nu/\nu_{0})^{6/5}$ $\mu$as = 1.8 $\mu$as at 5.5 GHz (Granot & van der Horst, 2014). This lower limit on the angular size can then be converted into constraints on the Lorentz factor as $\Gamma\gtrsim\theta_{d}D_{A}/c\Delta t_{rf}\approx 1.5$ at 10 d (rest-frame), where the angular diameter distance $D_{A}=D_{L}/(1+z)^{2}=4.8\times 10^{27}$ cm, $D_{L}$ is the luminosity distance, $c$ the speed of light, and $\Delta t_{rf}$ is the time since the burst in the rest-frame. Considering that the Lorentz factor of an expanding fireball decreases in time as $\Gamma\propto t^{-3/8}$ in an uniform density medium (Piran, 1999), we derive a lower limit on the initial Lorentz factor of $\Gamma\gtrsim 3.8$ at 1.5 d (observer’s frame), since at this time the optical and X-ray afterglow are seen to fade in time (Busmann et al., 2025), thus indicating that the blastwave is already in its deceleration phase. Our limit is rather conservative and the resulting conclusion of a mildly relativistic outflow does not strongly depend on the specific choice of input parameters. For instance, we verified that using the default settings of the NE2001 model, our limit would increase by a factor of $\approx$2.
In the case of refractive scintillation, the scattering disk is much larger than the first Fresnel zone, and this results in ${\theta_{r}=\theta_{\mathrm{F0}}(\nu/\nu_{0})^{-11/5}=2.3\,\mu\mathrm{as}}$ at 5.5 GHz (Granot & van der Horst, 2014). Large amplitude fluctuations are expected at 5.5 GHz, with modulation index $m_{r}=(\nu/\nu_{0})^{17/30}\sim 0.95$ (Granot & van der Horst, 2014), and at 2.1 GHz with $m_{r}\sim 0.6$. This effect is observed up to $\approx 25$ d at 5.5 GHz and up to $\approx 32$ d at 2.1 GHz. It explains the large discrepancy between the measured fluxes at different frequencies, but is likely quenched afterwards.
@@
Therefore, the observed variability of the radio counterpart consistently points to an outflow that is at least mildly relativistic, with Lorentz factor $\Gamma\approx$4. Unfortunately, without a clear identification of the deceleration time (e.g., Molinari et al., 2007), we cannot exclude the presence of an ultrarelativistic outflow at earlier times.
@@ 4.2.1 Variability function
The above estimates depend on the adopted ISM model, in particular the scattering measure. An alternative way to estimate this quantity directly from the data is by examining the variability timescale using intensity structure functions (Simonetti et al., 1985; Hughes et al., 1992; Ciaramella et al., 2004; Chandra et al., 2008),
@@
which is smaller than the NE2001 value. As the critical threshold depends on the SM as $\theta_{d}\propto SM^{3/5}\approx$0.9 $\mu$as (Cordes & Lazio, 2002), the resulting constraint on the Lorentz factor would only slightly weaken to $\Gamma\,\approx$2 at 1.5 d, which is still mildly relativistic.
@@ 4.3 Brightness Temperature
The brightness temperature $T_{b}$ provides an estimate of the intensity of radiation emitted by a source, assuming it behaves like a blackbody. Its value can be derived from early-time radio observations to aid constraining the relativistic nature of the outflow (Anderson et al., 2014).
For EP241021a, we use the first ATCA detection at $\sim$8 days for calculating $T_{b}$ as (Longair, 2011):
@@
where $S_{\nu}$ is the flux in Jy, $\nu$ is the frequency in Hz, $t$ is the time since the trigger in seconds, $z$ is the redshift. Since the maximum possible emitting region size is constrained by the speed of light ($ct$), comparing the observed $T_{b}$ with the inverse-Compton limit ($T_{B}\approx 10^{12}$ K) allows us to estimate the minimum bulk Lorentz factor ($\Gamma$) of the emitting material. This relationship is given by (Galama et al., 1999):
@@ 5 Multi-wavelength Afterglow Modeling
In the previous sections, we demonstrated that radio observations point to an expanding outflow with at least mildly relativistic velocity. The interaction of this fireball with the surrounding medium is known to produce broadband synchrotron radiation (Sari et al., 1998; Mészáros et al., 1999; Granot & Sari, 2002), known as afterglow, whose brightness and temporal evolution depend, among other factors, on the outflow structure, its geometry, and the observer’s viewing angle (Rossi et al., 1993; Sari et al., 1999; Troja et al., 2019; Ryan et al., 2020, 2024). This afterglow emission is generally described with a power-law form: $F_{\nu}\propto\nu^{-\beta}t^{-\alpha}$, where $\alpha$ and $\beta$ obey to standard closure relations (Zhang & Mészáros, 2004). We use this empirical model to derive basic physical constraints from the afterglow observations.
@@
  • the spectral index between 5.5 and 9 GHz is seen to vary. At early times, its value might be affected by ISS, which would explain its rapid variations. However, it displays a consistent trend of increasing from $\beta_{1}\sim-1/3$ to $\beta_{2}\sim 0.75$. This is consistent with the characteristic synchrotron frequency $\nu_{m}$ crossing this frequency range and progressively moving below 5.5 GHz after $T_{0}+70$ days.
@@
We use the best sampled light curve at 5.5 GHz to characterize the afterglow temporal evolution. We use a smoothly broken power law to fit the data, as described in Ricci et al. (2025). The flux is seen to rise with index $\alpha_{1}$=0.49$\pm$0.08 (68% confidence level), peak at around 40 d (rest-frame), then decay with slope $\alpha_{2}$= 2.6$\pm$0.4 (68% c. l.). The early rise, although likely affected by ISS, is consistent with the expected behavior for $\nu_{r}\lesssim\nu_{m}$, as also suggested by the spectral index $\beta_{1}\approx$-1/3. The steep late-time decay helps further constraining models as it disfavors quasi-spherical outflows, such as mildly relativistic cocoon or choked jet models (e.g., Troja et al., 2019), and points to a collimated outflow.
@@
Our empirical analysis shows that the radio emission in the range 2-17 GHz is broadly consistent with the evolution of optically thin synchrotron radiation from a population of shock-accelerated electrons. This allows us to place an upper limit on the self-absorption frequency $\nu_{sa}\lesssim$5 GHz, which in turn provides an independent constraint on the bulk Lorentz factor, $\Gamma\approx$4 (Barniol Duran et al., 2013). In contrast, the optical and near-infrared counterpart of EP241021a display a much more complex behavior, with a prominent bump at $\approx$7 d followed by a shallow long-lasting decay (Busmann et al., 2025). Non-standard evolution is also observed at X-ray energies (Sec. 3.3). Therefore, we exclude the late-time ($\gtrsim$7 d) optical and X-ray data from our modeling and test whether a basic afterglow model can consistently describe the radio dataset and the early optical and X-ray observations.
@@
Figure 6: Broadband modeling results. Best-fit afterglow light curves for an on-axis jet (model Top-Hat) compared with observations at multiple frequencies. Optical/nIR observations in r, z and J band are from Busmann et al. (2025) and radio observations at 15.5 GHz from Carotenuto et al. (AMI-LA; 2024). Downward arrows indicate $3\sigma$ upper limits. Empty symbols show data not used for model fitting.
@@
To derive physical constraints on the explosion, we model the afterglow using the open-source Python package Afterglowpy (Ryan et al., 2020, 2024), which implements the single-shell approximation (van Eerten et al., 2010; van Eerten, 2018). In this framework, the ejecta and forward shock are treated as a single radially uniform fluid element. The model incorporates prescriptions for jet spreading and uses an equation of state that smoothly transitions between the ultra-relativistic and non-relativistic regimes (van Eerten, 2013; Nava et al., 2013). Afterglowpy supports various jet structures, including Gaussian and power-law jets, by decomposing the jet into multiple top-hat segments. The dynamical evolution of each segment is governed by coupled ordinary differential equations that describe the shock radius, dimensionless four-velocity, and evolving jet opening angle. These equations are integrated over time to generate synthetic afterglow light curves.
@@
$\log\theta_{c}$ (rad) (-3, -0.5) $-0.8^{+0.1}_{-0.1}$ $-1.2^{+0.2}_{-0.1}$ $-0.9^{+0.02}_{-0.02}$
$\log\theta_{\rm obs}$ (rad) (-2, -0.5) $<-1.5$ $-0.9^{+0.2}_{-0.2}$ $-0.5$ (Fixed)
$\log n$ (cm-3) (-6, 2) $-3.7^{+0.7}_{-0.6}$ $-2.6^{+1.1}_{-0.9}$ $-0.3^{+0.3}_{-0.2}$
@@
$\log\epsilon_{e}$ (-6, -0.5) $-1.2^{+0.2}_{-0.1}$ $-1.1^{+0.2}_{-0.2}$ $-0.62^{+0.08}_{-0.09}$
$\log\epsilon_{B}$ (-6, -0.5) $>-2.2$ $-1.6^{+0.6}_{-0.8}$ $-2.6^{+0.7}_{-0.7}$
$\log\Gamma_{0}$ (0.1, 5) $1.6^{+0.1}_{-0.1}$ $>2.1$ $>2.5$
@@
We adopt a simple uniform (top-hat) jet profile to fit the afterglow data (see Fig.6) and treat the initial Lorentz factor $\Gamma_{0}$ as a free parameter [^3] A finite initial Lorentz factor breaks the assumption of instantaneous deceleration and necessitates disabling jet spreading, as standard spreading prescriptions are invalid under finite-$\Gamma_{0}$ conditions (Ryan et al., 2020). Accordingly, we set spread=False. For further discussion on the implications of finite initial Lorentz factors in Afterglowpy modeling, see Kaur et al. (2024).. The other free parameters include the isotropic-equivalent kinetic energy $E_{K,\mathrm{iso}}$, ambient density $n$, viewing angle $\theta_{\mathrm{obs}}$, jet core half-opening angle $\theta_{c}$, electron power-law index $p$, and the microphysical energy fractions in electrons ($\varepsilon_{e}$) and magnetic field ($\varepsilon_{B}$).
@@
The posterior distributions of these parameters are summarized in Table 3 and visualized in Fig. 7. This model favors an on-axis jet ($\theta_{\rm obs}<\theta_{\rm c}$) with moderate initial Lorentz factor $\Gamma_{0}\sim$30-50. As our line of sight intercepts the jet core, we are not particularly sensitive to the jet structure, and the simple top-hat profile provides an adequate, although not unique, description. Adopting the median values of $E_{K,\mathrm{iso}}\sim 1.3\times 10^{53}\,\mathrm{erg}$ and the jet half opening angle $\theta_{j}\approx 0.16$ rad ($\sim 9.2^{\circ}$; see Fig. 7), the beaming corrected kinetic energy is $E_{K}=(1-\cos\theta_{j})E_{k,\mathrm{iso}}\approx 1.6\times 10^{51}\,\text{ erg}$, consistent with relativistic TDEs and typical GRB energy scales (Wang et al., 2018; Beniamini et al., 2023).
@@
We also tested the model of a structured jet with a Gaussian angular profile. This model provides an adequate description of the data, but points to a highly relativistic outflow with $\Gamma$$>$130 seen off-axis ($\theta_{\rm obs}>\theta_{\rm c}$). Adopting the median values of $E_{K,\mathrm{iso}}\sim 5\times 10^{53}~{}\mathrm{erg}$ and a jet half-opening angle $\theta_{\rm c}\approx 0.06~{}\mathrm{rad}$ ($\sim 3.4^{\circ}$), the inferred beaming-corrected kinetic energy is $E_{K}\approx 10^{51}~{}\mathrm{erg}$. In this model, the jet core has a narrower opening angle than in the top-hat fit, but is surrounded by a wider angle structure from which emission is seen. Using this model, we can also reproduce the X-ray luminosity of the initial WXT detection (col. Gaussian+WXT). In this scenario, the early WXT flare would mark the afterglow onset rather than the time of the explosion.
@@
In all cases, the derived blastwave kinetic energy is much higher than the upper limit on $E_{\gamma,\rm iso}\lesssim 8\times 10^{50}$ erg (Sect. 1). This results in a radiative efficiency which is unusually low for classical GRBs, $\eta_{\gamma}=E_{\gamma,\rm iso}/(E_{\gamma,\rm iso}+E_{\rm K,iso})\lesssim$1%, although sporadically observed in some events (e.g., GRB190829A, Dichiara et al. 2022; and the candidate off-axis GRB150101B, Troja et al. 2018)
Figure 7: Broadband modelling results Corner plot showing the posterior distributions of the fitted model parameters using a uniform (top-hat) jet model.

### v2

@@ 3.4 X-ray Observations
Observations of EP241021a were conducted with the Follow-up X-ray Telescope (FXT) onboard Einstein Probe, starting $\approx$36.5 hours after the EP-WXT trigger (Y. Wang et al., 2024a). Follow-up observations from the Swift X-ray Telescope (XRT) were reported by M. Busmann et al. (2025), revealing a long-lasting X-ray counterpart. To determine whether the long-lived nature of the X-ray emission was possibly due to contamination from a nearby source, we analyzed public data from the Chandra X-ray Observatory contained in the Chandra Data Collection https://doi.org/10.25574/cdc.415 (catalog DOI: 10.25574/cdc.415), which observed EP241021a approximately 14 days after trigger (ObsID: 30598; PI: Jonker) for a total exposure of 9.9 ks. A single X-ray source is detected within the XRT localization at coordinates RA = 01h 55m 23.3984s, Dec = +05 56 17.498′′ (RA = 28.847, Dec = +5.938) with $90\%$ positional uncertainty of $0.52^{\prime\prime}$. Source counts were extracted using the CIAO software package (v4.16) [^3] https://cxc.harvard.edu/ciao/, employing a circular aperture with a radius of 1.1 arcseconds, corresponding to approximately 90% encircled energy of the Chandra point spread function (PSF). Aperture photometry was performed in the 0.5–8.0 keV band. The resulting spectrum was modeled using an absorbed power-law model (tbabs*powerlaw) in XSPEC, with the hydrogen column density ($N_{\mathrm{H}}$) fixed to the Galactic value reported in J. W. Hu et al. (2024b). We find a best-fit photon index of $\Gamma=1.3\pm 0.7$, and an unabsorbed model flux of $(6.1^{+1.4}_{-1.2})\times 10^{-14}$ erg cm-2 s-1 in the 0.3–10 keV band. This value is consistent with the nearly simultaneous XRT measurements (M. Busmann et al., 2025) and confirms that the late-time flattening is an intrinsic property of the source.
@@ 4 Constraints from radio observations
As mentioned in the Section 1, afterglows from compact radio sources are subject to ISS, which can produce significant flux modulations due to distorted wavefronts (M. A. Walker, 1998). Specifically, such flux variations can arise from diffractive and refractive scintillation, each sensitive to different angular scales and frequency bandwidths (R. Narayan, 1992; J. Goodman, 1997). By identifying the epochs when diffractive scintillation ceases and refractive scintillation becomes suppressed, we can place lower and upper bounds, respectively, on the angular size of the source. These size constraints, when combined with the angular diameter distance to the source, allow us to infer limits on the Lorentz factor of the outflow, as demonstrated in previous studies of GRBs and other relativistic transients (e.g., D. A. Frail et al., 2000; S. B. Cenko et al., 2008; L.-L. Zhang et al., 2023). Motivated by this, in Section 4.1, we present evidence for ISS during the radio monitoring of EP241021a. In Section 4.2, we use the ISS behavior to constrain the source’s angular size and Lorentz factor. We calculate the brightness temperature in Section 4.3 using early-time radio data.
@@ 4.1 Evidence for ISS
The strength and timescale of ISS effects depend on both the observing frequency and the angular size of the source (e.g., Y. Liu et al., 2022a). ISS manifests on two distinct timescales: diffractive ISS (DISS), which induces large-amplitude and rapid fluctuations over minutes due to small-scale interference patterns; and refractive ISS (RISS), which causes slower variability over weeks to months as the signal traverses larger-scale density structures in the scattering screen (M. A. Walker, 1998; W. Lewandowski et al., 2011; J. Granot & A. J. van der Horst, 2014; Y. Liu et al., 2022b). These scintillation effects can produce substantial variations in both flux and spectral shape, especially during the early stages of an afterglow. Initially, when the GRB jet is highly-collimated, DISS dominates the variability; as the emitting region expands, DISS becomes suppressed, and RISS emerges as the dominant contributor (D. A. Frail et al., 1997). Consequently, different radio facilities, operating at various frequencies and timescales, might detect variability on distinct timescales.
@@
Assuming that scattering occurs in a thin screen along the line of sight, the screen is located at a distance $D_{\text{scr}}\sim(h_{z}/2)\sin^{-1}|b|\approx 0.6$ kpc from the observer (R. J. Reynolds, 1989; T. J. Galama et al., 2003), where $h_{z}$ is the scale height of the ionised gas layer ($\sim 1$ kpc) and b=53.47b=-53.47 is the Galactic latitude in degrees. We use the NE2001 model of the Galactic distribution of free electrons (J. M. Cordes & T. J. W. Lazio, 2002; M. A. Walker, 1998) to derive the scattering measure (SM) of $1.3\times 10^{-4}\,\text{kpc}\,\text{m}^{-20/3}$ and a transition frequency of $\nu_{0}\sim 6$ GHz. Below this frequency, density irregularities in the ISM cause strong phase perturbations leading to significant scattering, whereas above this frequency, scattering effects become weaker.
The decorrelation bandwidth ($\Delta\nu_{\rm dc}$) for diffractive scintillation represents the frequency range over which the intensity variations caused by diffraction remain correlated (J. Goodman, 1997):
@@ 4.2 Angular size measurements and Lorentz factor
[…]
In the case of refractive scintillation, the scattering disk is much larger than the first Fresnel zone, and this results in $\theta_{r}=\theta_{\mathrm{F0}}(\nu/\nu_{0})^{-11/5}=2.3\,\mu\mathrm{as}$} at 5.5 GHz (J. Granot & A. J. van der Horst, 2014). Large amplitude fluctuations are expected at 5.5 GHz, with modulation index $m_{r}=(\nu/\nu_{0})^{17/30}\sim 0.95$ (J. Granot & A. J. van der Horst, 2014), and at 2.1 GHz with $m_{r}\sim 0.6$. This effect is observed up to $\approx 25$ d at 5.5 GHz and up to $\approx 32$ d at 2.1 GHz. It explains the large discrepancy between the measured fluxes at different frequencies, but is likely quenched afterwards.
@@
Therefore, the observed variability of the radio counterpart consistently points to an outflow that is at least mildly relativistic, with Lorentz factor $\Gamma\approx$4. Unfortunately, without a clear identification of the deceleration time (e.g., E. Molinari et al., 2007), we cannot exclude the presence of an ultrarelativistic outflow at earlier times.
@@ 4.2.1 Variability function
The above estimates depend on the adopted ISM model, in particular the scattering measure. An alternative way to estimate this quantity directly from the data is by examining the variability timescale using intensity structure functions (J. H. Simonetti et al., 1985; P. A. Hughes et al., 1992; A. Ciaramella et al., 2004; P. Chandra et al., 2008),
@@
[…]
@@ 4.3 Brightness Temperature
The brightness temperature $T_{b}$ provides an estimate of the intensity of radiation emitted by a source, assuming it behaves like a blackbody. Its value can be derived from early-time radio observations to aid constraining the relativistic nature of the outflow (G. E. Anderson et al., 2014).
For EP241021a, we use the first ATCA detection at $\sim$8 days for calculating $T_{b}$ as (M. S. Longair, 2011):
@@
where $S_{\nu}$ is the flux in Jy, $\nu$ is the frequency in Hz, $t$ is the time since the trigger in seconds, $z$ is the redshift. Since the maximum possible emitting region size is constrained by the speed of light ($ct$), comparing the observed $T_{b}$ with the inverse-Compton limit ($T_{B}\approx 10^{12}$ K) allows us to estimate the minimum bulk Lorentz factor ($\Gamma$) of the emitting material. This relationship is given by (T. J. Galama et al., 1999):
@@ 5 Multi-wavelength Afterglow Modeling
In the previous sections, we demonstrated that radio observations point to an expanding outflow with at least mildly relativistic velocity. The interaction of this fireball with the surrounding medium is known to produce broadband synchrotron radiation (R. Sari et al., 1998; P. Mészáros et al., 1999; J. Granot & R. Sari, 2002), known as afterglow, whose brightness and temporal evolution depend, among other factors, on the outflow structure, its geometry, and the observer’s viewing angle (P. Rossi et al., 1993; R. Sari et al., 1999; E. Troja et al., 2019; G. Ryan et al., 2020, 2024). This afterglow emission is generally described with a power-law form: $F_{\nu}\propto\nu^{-\beta}t^{-\alpha}$, where $\alpha$ and $\beta$ obey to standard closure relations (R. Sari et al., 1999; B. Zhang & P. Mészáros, 2004). We use this empirical model to derive basic physical constraints from the afterglow observations.
@@
  • the spectral index between 5.5 and 9 GHz is seen to vary. At early times, its value might be affected by ISS, which would explain its rapid variations. However, it displays a consistent trend of increasing from $\beta_{1}\sim-1/3$ to $\beta_{2}\sim 0.75$. This is consistent with the characteristic synchrotron frequency $\nu_{m}$ crossing this frequency range and progressively moving below 5.5 GHz after $T_{0}+70$ days, as expected for an external forward shock expanding into a uniform medium.
@@
We use the best-sampled radio light curve at 5.5 GHz, including our 6 GHz observations, to characterize the afterglow temporal evolution. A smoothly broken power-law fit (R. Ricci et al., 2025) yields a rising slope of $\alpha_{1}=0.7\pm 0.08$ (68% confidence level), peaking at $\sim$40 d (rest frame), followed by a decay with slope $\alpha_{2}=1.9\pm 0.1$ (68% c.l.).
@@
The early rise, although likely affected by ISS, is consistent with the expected behavior for $\nu_{r}\lesssim\nu_{m}$, as also suggested by the spectral index $\beta_{1}\approx$-1/3. The steep late-time decay helps further constraining models as it disfavors quasi-spherical outflows, such as mildly relativistic cocoon or choked jet models (e.g., E. Troja et al., 2019), and points to a collimated outflow.
@@
Our empirical analysis shows that the radio emission in the range 2-17 GHz is broadly consistent with the evolution of optically thin synchrotron radiation from a population of shock-accelerated electrons. This allows us to place an upper limit on the self-absorption frequency $\nu_{sa}\lesssim$5 GHz, which in turn provides an independent constraint on the bulk Lorentz factor, $\Gamma\approx$4 (R. Barniol Duran et al., 2013). In contrast, the optical and near-infrared counterpart of EP241021a display a much more complex behavior, with a prominent bump at $\approx$7 d followed by a shallow long-lasting decay (M. Busmann et al., 2025). Non-standard evolution is also observed at X-ray energies (Sec. 3.4). Therefore, we exclude the late-time ($\gtrsim$7 d) optical and X-ray data from our modeling and test whether a basic afterglow model can consistently describe the radio dataset and the early optical and X-ray observations.
@@
Figure 6: Broadband modeling results. Best-fit afterglow light curves for an on-axis jet (model Top-Hat) compared with observations at multiple frequencies. The X-ray (Swift/XRT; 1 keV) and optical and near-infrared observations in the r, z and J filters are from M. Busmann et al. (2025). Radio observations at 15.5 GHz from AMI-LA (F. Carotenuto et al., 2024). Downward arrows indicate $3\sigma$ upper limits. Empty symbols show data not used for model fitting. A scaling factor has been applied to different frequencies for plotting purposes.
Figure 7: Broadband modeling results. Corner plot showing the posterior distributions of the fitted model parameters using a uniform (top-hat) jet model.
@@
To derive physical constraints on the explosion, we model the afterglow using the open-source Python package Afterglowpy v. 0.8.0 (G. Ryan et al., 2020, 2024), which implements the single-shell approximation (H. van Eerten et al., 2010; H. van Eerten, 2018). In this framework, the ejecta and forward shock are treated as a single radially uniform fluid element. The model incorporates prescriptions for jet spreading and uses an equation of state that smoothly transitions between the ultra-relativistic and non-relativistic regimes (H. van Eerten, 2013; L. Nava et al., 2013). Afterglowpy supports various jet structures, including Gaussian and power-law jets, by decomposing the jet into multiple top-hat segments. The dynamical evolution of each segment is governed by coupled ordinary differential equations that describe the shock radius, dimensionless four-velocity, and evolving jet opening angle. These equations are integrated over time to generate synthetic afterglow light curves.
@@
$\log\theta_{c}$ (rad) (-3, -0.5) $0.8^{+0.1}_{-0.1}$ $1.2^{+0.2}_{-0.1}$ $0.9^{+0.02}_{-0.02}$
$\log\theta_{\rm obs}$ (rad) (-2, -0.5) <1.5<-1.5 $0.9^{+0.2}_{-0.2}$ 0.5-0.5 (Fixed)
$\log n$ (cm-3) (-6, 2) $3.7^{+0.7}_{-0.6}$ $2.6^{+1.1}_{-0.9}$ $0.3^{+0.3}_{-0.2}$
@@
$\log\epsilon_{e}$ (-6, -0.5) $2^{+0.2}_{-0.1} 1.1$$1^{+0.2}_{-0.2}$ $0.62^{+0.08}_{-0.09}$
$\log\epsilon_{B}$ (-6, -0.5) >2.2>-2.2 $1.6^{+0.6}_{-0.8}$ $2.6^{+0.7}_{-0.7}$
$\log\Gamma_{0}$ (0.1, 5) $1.6^{+0.1}_{-0.1}$ >2.1>2.1 >2.5>2.5
@@
We adopt a simple uniform (top-hat) jet profile to fit the afterglow data (see Fig.10) and treat the initial Lorentz factor $\Gamma_{0}$ as a free parameter [^4] A finite initial Lorentz factor breaks the assumption of instantaneous deceleration and necessitates disabling jet spreading, as standard spreading prescriptions are invalid under finite-$\Gamma_{0}$ conditions (G. Ryan et al., 2020). Accordingly, we set spread=False. For further discussion on the implications of finite initial Lorentz factors in Afterglowpy modeling, see R. Kaur et al. (2024).. The other free parameters include the isotropic-equivalent kinetic energy $E_{K,\mathrm{iso}}$, ambient density $n$, viewing angle $\theta_{\mathrm{obs}}$, jet core half-opening angle $\theta_{c}$, electron power-law index $p$, and the microphysical energy fractions in electrons ($\varepsilon_{e}$) and magnetic field ($\varepsilon_{B}$).
@@
The posterior distributions of these parameters are summarized in Table 3 and visualized in Fig. 10. This model favors an on-axis jet ($\theta_{\rm obs}<\theta_{\rm c}$) with moderate initial Lorentz factor $\Gamma_{0}\sim$30-50. As our line of sight intercepts the jet core, we are not particularly sensitive to the jet structure, and the simple top-hat profile provides an adequate, although not unique, description. Adopting the median values of $E_{K,\mathrm{iso}}\sim 1.3\times 10^{53}\,\mathrm{erg}$ and the jet half opening angle $\theta_{c}\approx 0.16$ rad ($\sim 9^{\circ}$; see Fig. 10), the beaming corrected kinetic energy is $E_{K}=(1-\cos\theta_{c})E_{k,\mathrm{iso}}\approx 1.6\times 10^{51}\,\text{ erg}$, consistent with relativistic TDEs and typical GRB energy scales (X.-G. Wang et al., 2018; P. Beniamini et al., 2023).
@@
We also tested the model of a structured jet with a Gaussian angular profile. This model provides an adequate description of the data, but points to a highly relativistic outflow with $\Gamma$>>130 seen off-axis ($\theta_{\rm obs}>\theta_{\rm c}$). Adopting the median values of $E_{K,\mathrm{iso}}\sim 5\times 10^{53}~\mathrm{erg}$ and a jet half-opening angle $\theta_{\rm c}\approx 0.06~\mathrm{rad}$ ($\sim 3.4^{\circ}$), the inferred beaming-corrected kinetic energy is $E_{K}\approx 10^{51}~\mathrm{erg}$. In this model, the jet core has a narrower opening angle than in the top-hat fit, but is surrounded by a wider angle structure from which emission is seen. Using this model, we can also reproduce the X-ray luminosity of the initial WXT detection (col. Gaussian+WXT). In this scenario, the early WXT flare would mark the afterglow onset rather than the time of the explosion.
@@
In all cases, the derived blastwave kinetic energy is much higher than the upper limit on $E_{\gamma,\rm iso}\lesssim 8\times 10^{50}$ erg (Sect. 1). This results in a radiative efficiency which is unusually low for classical GRBs, $\eta_{\gamma}=E_{\gamma,\rm iso}/(E_{\gamma,\rm iso}+E_{\rm K,iso})\lesssim$1%, although sporadically observed in some events (e.g., GRB190829A, S. Dichiara et al. 2022; and the candidate off-axis GRB150101B, E. Troja et al. 2018)